To achieve maximum range for a given initial velocity what will be the angle of projection?
1
Expert's answer
2020-07-02T17:15:25-0400

The expression for the range of a body launched with initial velocity "v" directed "\theta" degrees above the horizontal is


"R=\frac{v^2\text{ sin}(2\theta)}{g}."

We can see that the range is at maximum when "\theta=45°".